### v3

@@ 4 Unified Theoretical Analysis
In this section, we introduce the ideas in our analysis and present three important lemmas, all the missing proofs of which are deferred into Appendix B.
@@
where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right],\forall t\in\left[T\right]$ and $z^{t}\coloneqq\frac{v_{0}}{v_{t}}x+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right]$.
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}.$
@@ 5.2 New Assumption and Some Discussions
    Finite $p$-th moment: $\exists p\in\left(1,2\right)$, $\sigma\geq 0$ denoting the noise level such that $\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{p}\mid\mathcal{F}^{t-1}\right]\leq\sigma^{p}$.
@@ 5.3 General Convex Functions under Heavy-Tailed Noise
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{L\left[D_{\psi}(x^{*},x^{1})\right]^{2-\frac{2}{p}}\log T}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x^{*},x^{1})\right]^{1-\frac{1}{p}}\log T}{T^{1-\frac{1}{p}}}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{L\left[D_{\psi}(x^{*},x^{1})\right]^{2-\frac{2}{p}}(\log T)^{\frac{2}{p}-1}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x^{*},x^{1})\right]^{1-\frac{1}{p}}(\log T)^{\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
@@ 5.3.1 Optimal Rate under Heavy-Tailed Noise
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{L\left[D_{\psi}(x^{*},x^{1})\right]^{2-\frac{2}{p}}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x^{*},x^{1})\right]^{1-\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
@@ 6.2 New Assumption and Some Discussions
    Sub-Weibull noise: $\exists p\in\left(0,2\right)$, $\sigma\geq 0$ denoting the noise level such that $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi^{t}\right\|_{*}^{p}\right)\mid\mathcal{F}^{t-1}\right]\leq\exp\left(\lambda\sigma^{p}\right),\forall\lambda\in\left[0,\sigma^{-p}\right]$.
@@
Given a sigma algebra $\mathcal{F}$ and a random vector $Z\in\mathbb{R}^{d}$ that is $\mathcal{F}$-measurable, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{\mathbb{E}}\left[\xi\mid\mathcal{F}\right]=0$ and $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{p}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{p}\right),\forall\lambda\in\left[0,\sigma^{-p}\right]$ for $p\in\left[1,2\right)$, then
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\leq\begin{cases}\exp\left(6\sigma^{2}\left\|Z\right\|^{2}\right)&\text{if }\left\|Z\right\|\leq\frac{1}{2\sigma}\text{ almost surely}\\ \exp\left(6\sigma^{2}\left\|Z\right\|^{2}+2^{q-1}\sigma^{q}\left\|Z\right\|^{q}\right)&p\in\left(1,2\right),q\coloneqq\frac{p}{p-1}\end{cases}.$
@@ Appendix A Technical Lemmas
Given a sigma algebra $\mathcal{F}$, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in\left[0,\sigma^{-2}\right]$, then for any integer $k\geq 1$, we have
$\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2k}\mid\mathcal{F}\right]\leq\begin{cases}\sigma^{2}&k=1\\ e(k!)\sigma^{2k}&k\geq 2\end{cases}.$
@@
$\exp\left(\mathbb{\mathbb{E}}\left[\lambda\left\|\xi\right\|_{*}^{2}\mid\mathcal{F}\right]\right)\leq\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{2}\right)\Rightarrow\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2}\mid\mathcal{F}\right]\leq\sigma^{2}.$
@@
$\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2k}\mid\mathcal{F}\right]$ $=\mathbb{\mathbb{E}}\left[\int_{0}^{\infty}2kt^{2k-1}\mathds{1}\left[\left\|\xi\right\|_{*}\geq t\right]\mathrm{d}t\mid\mathcal{F}\right]=\int_{0}^{\infty}2kt^{2k-1}\mathbb{\mathbb{E}}\left[\mathds{1}\left[\left\|\xi\right\|_{*}\geq t\right]\mid\mathcal{F}\right]\mathrm{d}t$
$\leq\int_{0}^{\infty}2kt^{2k-1}\mathbb{\mathbb{E}}\left[\frac{\exp\left(\sigma^{-2}\left\|\xi\right\|_{*}^{2}\right)}{\exp(\sigma^{-2}t^{2})}\mid\mathcal{F}\right]\mathrm{d}t\overset{(a)}{\leq}\int_{0}^{\infty}2ekt^{2k-1}\exp(-\sigma^{-2}t^{2})\mathrm{d}t$
@@
where $(a)$ is by $\mathbb{\mathbb{E}}\left[\exp\left(\sigma^{-2}\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp(\sigma^{-2}\sigma^{2})=e$ and $(b)$ is by the change of variable $t=\sigma\sqrt{s}$. ∎
@@
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]$ $=\mathbb{\mathbb{E}}\left[1+\left\langle\xi,Z\right\rangle+\sum_{k=2}^{\infty}\frac{\left(\left\langle\xi,Z\right\rangle\right)^{k}}{k!}\mid\mathcal{F}\right]\leq\mathbb{\mathbb{E}}\left[\left\langle\xi,Z\right\rangle\mid\mathcal{F}\right]+\mathbb{\mathbb{E}}\left[1+\sum_{k=2}^{\infty}\frac{\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}}{k!}\mid\mathcal{F}\right]$
$\overset{(a)}{=}\mathbb{\mathbb{E}}\left[1+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}}{(2k)!}+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k+1}\left\|Z\right\|^{2k+1}}{(2k+1)!}\mid\mathcal{F}\right]$
$\overset{(b)}{\leq}\mathbb{\mathbb{E}}\left[1+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}}{(2k)!}+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}+\left\|\xi\right\|_{*}^{2k+2}\left\|Z\right\|^{2k+2}/4}{(2k+1)!}\mid\mathcal{F}\right]$
$=\mathbb{\mathbb{E}}\left[1+\frac{2\left\|\xi\right\|_{*}^{2}\left\|Z\right\|^{2}}{3}+\sum_{k=2}^{\infty}\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}\left(\frac{1}{4(2k-1)!}+\frac{1}{(2k)!}+\frac{1}{(2k+1)!}\right)\mid\mathcal{F}\right]$
$=\mathbb{\mathbb{E}}\left[1+\frac{2\left\|\xi\right\|_{*}^{2}\left\|Z\right\|^{2}}{3}+\sum_{k=2}^{\infty}\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}\frac{1+k/2+1/(2k+1)}{(2k)!}\mid\mathcal{F}\right]$
@@
where $(a)$ is by $\mathbb{\mathbb{E}}\left[\left\langle\xi,Z\right\rangle\mid\mathcal{F}\right]=\left\langle\mathbb{\mathbb{E}}\left[\xi\mid\mathcal{F}\right],Z\right\rangle=0$, $(b)$ holds due to AM-GM inequality, $(c)$ is by applying Lemma A.1 to $\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}\mid\mathcal{F}\right]=\left\|Z\right\|^{2k}\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{2k}\mid\mathcal{F}\right]$ and $(d)$ is by $2/3<1$ and
$\max_{k\geq 2,k\in\mathbb{\mathbb{N}}}\frac{e(1+k/2+1/(2k+1))}{\binom{2k}{k}}=\frac{e(1+1+1/5)}{6}<1.$
@@
Given a sigma algebra $\mathcal{F}$, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{p}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{p}\right),\forall\lambda\in\left[0,\sigma^{-p}\right]$ for some $p\in\left(0,2\right)$, then for any $k\geq 1$ (not necessarily an integer), we have
$\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{k}\mid\mathcal{F}\right]\leq\begin{cases}e(k!)\sigma^{k}&k\in\mathbb{\mathbb{N}}\text{ and }p\in\left[1,2\right)\\ e\left(\frac{k}{p}\right)^{\frac{k}{p}}\sigma^{k}&k\geq p\end{cases}.$
@@
$\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{k}\mid\mathcal{F}\right]\leq e\sigma^{k}\Gamma\left(\frac{k}{p}+1\right).$
@@
where the last step is by $\Gamma\left(\frac{k}{p}-\left\lfloor\frac{k}{p}\right\rfloor+1\right)\leq 1$ due to $1\leq\frac{k}{p}-\left\lfloor\frac{k}{p}\right\rfloor+1<2$. Finally, when $p\in\left[1,2\right)$ and $k\in\mathbb{\mathbb{N}}$, we can bound
@@
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]$ $=\mathbb{\mathbb{E}}\left[1+\left\langle\xi,Z\right\rangle+\sum_{k=2}^{\infty}\frac{\left(\left\langle\xi,Z\right\rangle\right)^{k}}{k!}\mid\mathcal{F}\right]\overset{(a)}{\leq}\mathbb{\mathbb{E}}\left[1+\sum_{k=2}^{\infty}\frac{\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}}{k!}\mid\mathcal{F}\right]$
@@
where $(a)$ is due to $\mathbb{\mathbb{E}}\left[\left\langle\xi,Z\right\rangle\mid\mathcal{F}\right]=\left\langle\mathbb{\mathbb{E}}\left[\xi\mid\mathcal{F}\right],Z\right\rangle=0$ and $\left(\left\langle\xi,Z\right\rangle\right)^{k}\leq\left(\left|\left\langle\xi,Z\right\rangle\right|\right)^{k}\leq\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}$, and $(b)$ is by applying Lemma A.2 to $\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}\mid\mathcal{F}\right]=\left\|Z\right\|^{k}\mathbb{\mathbb{E}}\left[\left\|\xi\right\|_{*}^{k}\mid\mathcal{F}\right]$.
@@
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\leq\mathbb{\mathbb{E}}\left[\exp\left(\frac{\left\|\xi\right\|_{*}^{p}}{2p\sigma^{p}}+\frac{2^{q/p}\sigma^{q}\left\|Z\right\|^{q}}{q}\right)\mid\mathcal{F}\right]\leq\exp\left(\frac{1}{2p}+\frac{2^{q/p}\sigma^{q}\left\|Z\right\|^{q}}{q}\right),$ (4)
where the last step is due to $Z\in\mathcal{F}$ and $\mathbb{\mathbb{E}}\left[\exp\left(\frac{\left\|\xi\right\|_{*}^{p}}{2p\sigma^{p}}\right)\mid\mathcal{F}\right]\leq\exp\left(\frac{1}{2p}\right)$ since $\frac{1}{2p}\leq 1$. Now, let us define the event $A\coloneqq\left\{\left\|Z\right\|\leq\frac{1}{2\sigma}\right\}$ and bound
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]$ $=\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\left(\mathds{1}\left[A\right]+\mathds{1}\left[A^{c}\right]\right)$

### v4

@@ 4 Unified Theoretical Analysis
In this section, we introduce the ideas in our analysis and present three important lemmas, all the missing proofs of which are deferred to Appendix B.
@@
where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{E}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right],\forall t\in\left[T\right]$ and $z^{t}\coloneqq\frac{v_{0}}{v_{t}}x+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right]$.
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}.$
@@ 5.2 New Assumption and Some Discussions
    Finite $p$-th moment: $\exists p\in\left(1,2\right)$, $\sigma\geq 0$ denoting the noise level such that $\mathbb{E}\left[\left\|\xi^{t}\right\|_{*}^{p}\right]\leq\sigma^{p}$.
@@ 5.3 General Convex Functions under Heavy-Tailed Noise
$\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{L\left[D_{\psi}(x^{*},x^{1})\right]^{2-\frac{2}{p}}\log T}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x^{*},x^{1})\right]^{1-\frac{1}{p}}\log T}{T^{1-\frac{1}{p}}}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{L\left[D_{\psi}(x^{*},x^{1})\right]^{2-\frac{2}{p}}(\log T)^{\frac{2}{p}-1}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x^{*},x^{1})\right]^{1-\frac{1}{p}}(\log T)^{\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
@@ 5.3.1 Optimal Rate under Heavy-Tailed Noise
$\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{L\left[D_{\psi}(x^{*},x^{1})\right]^{2-\frac{2}{p}}}{T^{2-\frac{2}{p}}}+\frac{(M+\sigma)\left[D_{\psi}(x^{*},x^{1})\right]^{1-\frac{1}{p}}}{T^{1-\frac{1}{p}}}\right).$
@@ 6.2 New Assumption and Some Discussions
    Sub-Weibull noise: $\exists p\in\left(0,2\right)$, $\sigma\geq 0$ denoting the noise level such that $\mathbb{E}\left[\exp\left(\lambda\left\|\xi^{t}\right\|_{*}^{p}\right)\mid\mathcal{F}^{t-1}\right]\leq\exp\left(\lambda\sigma^{p}\right),\forall\lambda\in\left[0,\sigma^{-p}\right]$.
@@
Given a sigma algebra $\mathcal{F}$ and a random vector $Z\in\mathbb{R}^{d}$ that is $\mathcal{F}$-measurable, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{E}\left[\xi\mid\mathcal{F}\right]=0$ and $\mathbb{E}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{p}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{p}\right),\forall\lambda\in\left[0,\sigma^{-p}\right]$ for $p\in\left[1,2\right)$, then
$\mathbb{E}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\leq\begin{cases}\exp\left(6\sigma^{2}\left\|Z\right\|^{2}\right)&\text{if }\left\|Z\right\|\leq\frac{1}{2\sigma}\text{ almost surely}\\ \exp\left(6\sigma^{2}\left\|Z\right\|^{2}+2^{q-1}\sigma^{q}\left\|Z\right\|^{q}\right)&p\in\left(1,2\right),q\coloneqq\frac{p}{p-1}\end{cases}.$
@@ Appendix A Technical Lemmas
Given a sigma algebra $\mathcal{F}$, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{E}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in\left[0,\sigma^{-2}\right]$, then for any integer $k\geq 1$, we have
$\mathbb{E}\left[\left\|\xi\right\|_{*}^{2k}\mid\mathcal{F}\right]\leq\begin{cases}\sigma^{2}&k=1\\ e(k!)\sigma^{2k}&k\geq 2\end{cases}.$
@@
$\exp\left(\mathbb{E}\left[\lambda\left\|\xi\right\|_{*}^{2}\mid\mathcal{F}\right]\right)\leq\mathbb{E}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{2}\right)\Rightarrow\mathbb{E}\left[\left\|\xi\right\|_{*}^{2}\mid\mathcal{F}\right]\leq\sigma^{2}.$
@@
$\mathbb{E}\left[\left\|\xi\right\|_{*}^{2k}\mid\mathcal{F}\right]$ $=\mathbb{E}\left[\int_{0}^{\infty}2kt^{2k-1}\mathds{1}\left[\left\|\xi\right\|_{*}\geq t\right]\mathrm{d}t\mid\mathcal{F}\right]=\int_{0}^{\infty}2kt^{2k-1}\mathbb{E}\left[\mathds{1}\left[\left\|\xi\right\|_{*}\geq t\right]\mid\mathcal{F}\right]\mathrm{d}t$
$\leq\int_{0}^{\infty}2kt^{2k-1}\mathbb{E}\left[\frac{\exp\left(\sigma^{-2}\left\|\xi\right\|_{*}^{2}\right)}{\exp(\sigma^{-2}t^{2})}\mid\mathcal{F}\right]\mathrm{d}t\overset{(a)}{\leq}\int_{0}^{\infty}2ekt^{2k-1}\exp(-\sigma^{-2}t^{2})\mathrm{d}t$
@@
where $(a)$ is by $\mathbb{E}\left[\exp\left(\sigma^{-2}\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp(\sigma^{-2}\sigma^{2})=e$ and $(b)$ is by the change of variable $t=\sigma\sqrt{s}$. ∎
@@
$\mathbb{E}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]$ $=\mathbb{E}\left[1+\left\langle\xi,Z\right\rangle+\sum_{k=2}^{\infty}\frac{\left(\left\langle\xi,Z\right\rangle\right)^{k}}{k!}\mid\mathcal{F}\right]\leq\mathbb{E}\left[\left\langle\xi,Z\right\rangle\mid\mathcal{F}\right]+\mathbb{E}\left[1+\sum_{k=2}^{\infty}\frac{\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}}{k!}\mid\mathcal{F}\right]$
$\overset{(a)}{=}\mathbb{E}\left[1+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}}{(2k)!}+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k+1}\left\|Z\right\|^{2k+1}}{(2k+1)!}\mid\mathcal{F}\right]$
$\overset{(b)}{\leq}\mathbb{E}\left[1+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}}{(2k)!}+\sum_{k=1}^{\infty}\frac{\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}+\left\|\xi\right\|_{*}^{2k+2}\left\|Z\right\|^{2k+2}/4}{(2k+1)!}\mid\mathcal{F}\right]$
$=\mathbb{E}\left[1+\frac{2\left\|\xi\right\|_{*}^{2}\left\|Z\right\|^{2}}{3}+\sum_{k=2}^{\infty}\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}\left(\frac{1}{4(2k-1)!}+\frac{1}{(2k)!}+\frac{1}{(2k+1)!}\right)\mid\mathcal{F}\right]$
$=\mathbb{E}\left[1+\frac{2\left\|\xi\right\|_{*}^{2}\left\|Z\right\|^{2}}{3}+\sum_{k=2}^{\infty}\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}\frac{1+k/2+1/(2k+1)}{(2k)!}\mid\mathcal{F}\right]$
@@
where $(a)$ is by $\mathbb{E}\left[\left\langle\xi,Z\right\rangle\mid\mathcal{F}\right]=\left\langle\mathbb{E}\left[\xi\mid\mathcal{F}\right],Z\right\rangle=0$, $(b)$ holds due to AM-GM inequality, $(c)$ is by applying Lemma A.1 to $\mathbb{E}\left[\left\|\xi\right\|_{*}^{2k}\left\|Z\right\|^{2k}\mid\mathcal{F}\right]=\left\|Z\right\|^{2k}\mathbb{E}\left[\left\|\xi\right\|_{*}^{2k}\mid\mathcal{F}\right]$ and $(d)$ is by $2/3<1$ and
$\max_{k\geq 2,k\in\mathbb{N}}\frac{e(1+k/2+1/(2k+1))}{\binom{2k}{k}}=\frac{e(1+1+1/5)}{6}<1.$
@@
Given a sigma algebra $\mathcal{F}$, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{E}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{p}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{p}\right),\forall\lambda\in\left[0,\sigma^{-p}\right]$ for some $p\in\left(0,2\right)$, then for any $k\geq 1$ (not necessarily an integer), we have
$\mathbb{E}\left[\left\|\xi\right\|_{*}^{k}\mid\mathcal{F}\right]\leq\begin{cases}e(k!)\sigma^{k}&k\in\mathbb{N}\text{ and }p\in\left[1,2\right)\\ e\left(\frac{k}{p}\right)^{\frac{k}{p}}\sigma^{k}&k\geq p\end{cases}.$
@@
$\mathbb{E}\left[\left\|\xi\right\|_{*}^{k}\mid\mathcal{F}\right]\leq e\sigma^{k}\Gamma\left(\frac{k}{p}+1\right).$
@@
where the last step is by $\Gamma\left(\frac{k}{p}-\left\lfloor\frac{k}{p}\right\rfloor+1\right)\leq 1$ due to $1\leq\frac{k}{p}-\left\lfloor\frac{k}{p}\right\rfloor+1<2$. Finally, when $p\in\left[1,2\right)$ and $k\in\mathbb{N}$, we can bound
@@
$\mathbb{E}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]$ $=\mathbb{E}\left[1+\left\langle\xi,Z\right\rangle+\sum_{k=2}^{\infty}\frac{\left(\left\langle\xi,Z\right\rangle\right)^{k}}{k!}\mid\mathcal{F}\right]\overset{(a)}{\leq}\mathbb{E}\left[1+\sum_{k=2}^{\infty}\frac{\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}}{k!}\mid\mathcal{F}\right]$
@@
where $(a)$ is due to $\mathbb{E}\left[\left\langle\xi,Z\right\rangle\mid\mathcal{F}\right]=\left\langle\mathbb{E}\left[\xi\mid\mathcal{F}\right],Z\right\rangle=0$ and $\left(\left\langle\xi,Z\right\rangle\right)^{k}\leq\left(\left|\left\langle\xi,Z\right\rangle\right|\right)^{k}\leq\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}$, and $(b)$ is by applying Lemma A.2 to $\mathbb{E}\left[\left\|\xi\right\|_{*}^{k}\left\|Z\right\|^{k}\mid\mathcal{F}\right]=\left\|Z\right\|^{k}\mathbb{E}\left[\left\|\xi\right\|_{*}^{k}\mid\mathcal{F}\right]$.
@@
$\mathbb{E}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\leq\mathbb{E}\left[\exp\left(\frac{\left\|\xi\right\|_{*}^{p}}{2p\sigma^{p}}+\frac{2^{q/p}\sigma^{q}\left\|Z\right\|^{q}}{q}\right)\mid\mathcal{F}\right]\leq\exp\left(\frac{1}{2p}+\frac{2^{q/p}\sigma^{q}\left\|Z\right\|^{q}}{q}\right),$ (4)
where the last step is due to $Z\in\mathcal{F}$ and $\mathbb{E}\left[\exp\left(\frac{\left\|\xi\right\|_{*}^{p}}{2p\sigma^{p}}\right)\mid\mathcal{F}\right]\leq\exp\left(\frac{1}{2p}\right)$ since $\frac{1}{2p}\leq 1$. Now, let us define the event $A\coloneqq\left\{\left\|Z\right\|\leq\frac{1}{2\sigma}\right\}$ and bound
$\mathbb{E}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]$ $=\mathbb{E}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\left(\mathds{1}\left[A\right]+\mathds{1}\left[A^{c}\right]\right)$